Three numbers added together total 68. The second number is four more than the first, the third number is twice the first. find the numbers.
Answer by ankor@dixie-net.com(22740)

Just write a simple equation for each statement
:
Three numbers added together total 68.
x + y + z = 68
:
The second number is four more than the first,
y = x + 4
:
the third number is twice the first.
z = 2x
:
find the numbers.
x + y + z = 68
Subsitute
x + (x+4) + 2x = 68
x + x + 4 + 2x = 68
4x = 68 - 4
4x = 64
x = 64/4
x = 16
:
y = 16 + 4 = 20
z = 2(16) = 32
:
Check: 16 + 20 + 32 = 68